Find an equation for the line with the given properties.
Parallel to the line 9x+4y=57; containing the point(5,-1)

Answer by jim_thompson5910(35256)

9x+4y=C


9(5)+4(-1)=C


45-4=C


41=C


So the parallel equation is 9x+4y=41